OpenStudy (anonymous):

solve the equation -9(t-7)-(t+2)=7

OpenStudy (anonymous):

did you understand?

OpenStudy (anonymous):

now write...

OpenStudy (anonymous):

no im still stuck

OpenStudy (anonymous):

I think my answer is wrong I came up with 61

OpenStudy (anonymous):

-9(t-7)-(t+2)=7 now multiply -9t+63-t-2=7 -10t+61=7 -10t=7-61 -10t=-54 t=54/10 t=5.4

OpenStudy (anonymous):

thanks I see the part I was missing now

OpenStudy (anonymous):

this will be confused maximum with the sighns